The October 1 dome continues to grow at a relatively high rate, and its eastern face has become steeper, making dome collapses more likely. The morphology of the dome has changed slightly, to a more blocky texture. Several inferred small explosive signals have been recorded, along with some rockfalls. The rockfalls are becoming larger. There was a small swarm of mostly shallow volcano-tectonic earthquakes.
The volcano was generally very cloudy, with only limited views of the dome complex possible until 18 October. Visibility was excellent on 18 and 19 October.
At the beginning of the week, continued growth of the October 1 dome was observed, along with a change in the dome morphology. The dome has been growing since the beginning of October within the scar left by the explosive event of 17/18 September. The dome surface has become blocky, and the dome now has a more similar appearance to previous domes. There were two peaks, at the southwest and eastern sides of the dome on 14 October. A small, new, stubby spine was noted on 15 October, and had disappeared by the following day.
By 18 October, the dome had grown rapidly on its north side, and became 15 metres higher than the pre-Sept 17 dome complex at the northern edge of the scar. Lateral expansion continued towards the west, so that the dome covered most of the floor of the scar. The faces of the dome became generally more blocky. The eastern face became quite steep, with several unstable-looking blocks poised at the top of the slope. Small rockfalls from the eastern face were observed throughout the week whenever the cloud lifted. Views from Windy Hill, to the NNW of the crater, were possible on 19 October; however, no discernable change to the shape of the dome complex was seen from there.
There has been an increase in the number of reported ash clouds. Although these ash clouds are small, reaching no more than 2000 ft above the crater area, their occurrence suggests that the dome surface has become more unstable. Collapses of the eastern face of the October 1 dome are anticipated in the near future, although it is not thought that there is enough unstable material available yet to generate large pyroclastic flows.
Earthquake types: 13 to 19 October 1996
These earthquake counts are of events that triggered the short-period seismic network event recording system between 0000 to 2400 each day.
Volcano-tectonic earthquakes and rockfall signals have dominated the seismicity. The general level of activity remains low, continuing the trend established after the 17/18 September explosion. However, the size of the rockfalls has increased slightly, which is consistent with continued dome growth and the dome slopes becoming less stable.
Date VT LP Hybrid RF Tremor 13 OCT 96 5 2 0 5 low 14 OCT 96 6 0 0 5 moderate 15 OCT 96 4 5 0 5 low 16 OCT 96 4 4 0 8 moderate 17 OCT 96 11 3 0 9 moderate 18 OCT 96 32 4 0 9 low 19 OCT 96 6 0 0 7 low
The volcano-tectonic earthquakes have mostly been located beneath the crater at very shallow depths, and most of the events occurred during swarms from 2200 17 October to 0430 18 October, and 1300 18 October to 0130 19 October. Many of the locations have been above 1 km depth, and the seismic signals show extended long-period codas at some stations, consistent with very shallow hypocentres. There were two deeper earthquakes, one at about 6 km depth near Roaches Yard, and the other at about 15 km depth, 4 km to the east of the crater.
Several long-period earthquakes have been recorded. Often these events are immediately followed by rockfall signals, and they are interpreted as small explosions close to the surface of the dome which have caused collapses from the dome. The largest event, at 12:00 on 15 October, produced a small ash column.
The new broadband network is now installed and operating, although there is still a minor problem with the telephone link from the station at Long Ground. The seven other stations, five of which have three-component, broadband sensors, are transmitting 24-bit data to the Observatory successfully. The data are recorded continuously, and also processed with trigger algorithms to detect events of different types. Preliminary investigations show that the data are of high quality, and some short-duration, low-frequency, tremor-like signals have been detected that did not trigger the short-period network. The two networks will run in parallel, at least for the next few months.
EDM measurements were carried out on the eastern triangle between Whites, Long Ground and Castle Peak on 15, 18 and 19 October. On the 15th, the line between Long Ground and Whites could not be completed because of poor visibility. The Whites to Castle Peak line shortened by 2.5 cm during the period 10 to 19 October, and the Long Ground to Castle Peak line shortened by 3.4 cm from 10 to 19 October. These results indicate the eastward movement of Castle Peak of 2-3 mm per day continues, a the trend established in mid-July.
The southern triangle between Galways, O'Garros and Chances Peak was measured on 19 October. The Galways Castle Peak line lengthened by 3.8 cm from 6 October. This is the second successive lengthening of this line.
A GPS survey on the MVO Bignet which straddles the volcano was completed on 13 October. Preliminary processing of the results shows no significant changes of the line lengths since the network was last measured after the 17/18 September explosion.
The University of Puerto Rico GPS program of permanent GPS monitoring has been scaled down due to the destruction of one of the two stations during the 17/18 September explosion. Currently one site, at Reid's Hill to the southwest of the volcano, is permanently occupied. There are plans to establish another site at Whites Yard.
Dome volume measurements
During the excellent viewing conditions of 18 October a comprehensive dome survey was undertaken from the helicopter, using the GPS equipment to determine the location of the helicopter and laser-ranging binoculars to map the surface of the dome.
The October 1 dome had a volume of 4.7 x 106 m3 on 18 October, which gives an effusion rate of 3 m3/s between 1 and 18 October. The scar volume was estimated at 9 x 106 m3, approximately one-third of the volume of the dome complex before the explosion.
A limited bathymetric survey of the delta was carried out by boat on 16 October. The nose of the delta is steep, and reaches depths of 20 metres within a short distance of the shore line. Further surveys using sonar equipment are planned.
Gas and ash and rainwater measurements
The correlation spectrometer has returned to Montserrat after being repaired, which has allowed the SO2 flux to be estimated for the first time since 13 August. Surveys were carried out on 13, 15 and 17 October. The weather conditions were unsuitable on other days.
The average flux readings were 550, 650 and 1110 tonnes per day, respectively. The latter value is considered fairly high for a sulphur-poor volcano such as the Soufriere Hills, and is similar to readings obtained in late July and early August. These readings suggest that gas-rich magma is reaching the surface.
Results of gas analysis of the Galways Soufriere on 4 July 96 were received from the Volcano Observatory in Guadeloupe. There were no significant changes in gas composition compared with earlier samples. The gas is dominated by CO2 (60%) and H2S (39%).
The SO2 tubes at sites around Amersham and Plymouth are collected every two weeks and sent to the UK for analysis. No results have been obtained this week. No ash was sampled this week. Weekly collecting and analysis of rainwater continues.
Dr Paul Jackson, SRU
Chloe Harford, Bristol University, U.K.
Dave Stewart, BGS
Peter Day, BGS
Mr Lloyd Lynch, SRU
Prof Geoff Wadge, Reading University, U.K.